# Arrow-Debreu Meets Kyle: Price Discovery Across Derivatives

### v35

@@ 1 Introduction
A fundamental economic function of prices is to convey information. As informed agents trade to exploit their information about an asset’s expected payoff, that information becomes incorporated into the asset price. More broadly, agents may have information about aspects of an asset’s payoff beyond just its expected value—such as volatility—which can only be exploited through trading derivatives. This role of derivatives in providing targeted exposure to various aspects of an asset’s payoff distribution has been well understood since Ross (1976). Indeed, such derivatives trading strategies have long been standard practice, as reflected in widely used textbooks like Hull (2015). This implies that a broad range of information beyond an asset’s expected payoff is incorporated in derivatives prices.
Yet, the existing literature lacks a basic model articulating how derivative markets systematically incorporate diverse information signals into prices. This gap stems largely from the challenges of modeling the complex information dynamics across derivatives. We address this gap by developing a model that unifies the foundational frameworks of Arrow and Debreu (1954) and Kyle (1985). The model fully captures the nonlinearity of payoff distributions and the interdependence of trades and prices across derivatives. We obtain general characterizations of the informed demand, price impact, and information efficiency of prices across derivative markets.
In our model, an informed trader has private information about the probabilities of future states and trades state-contingent claims. In other words, the informed trader has information about the payoffs of Arrow-Debreu (AD) securities and trades in the AD markets. There is no restriction on the probability distribution across states or the type of information available to the informed trader.
Replicating derivatives—such as options—using AD securities yields alternative formulations of our model that account for the structural linkages across derivative instruments. In an options formulation, the informed trader has arbitrary private information about an underlying asset’s payoff distribution and trades options.
@@
The joint supply schedules for all AD securities form a pricing kernel that prices all derivatives simultaneously under asymmetric information. While derivatives pricing has been extensively studied from the hedging perspective since Black and Scholes (1973), we move beyond the limitations of complete-information hedging models and show that informed trading explains observed derivative price behaviors that the hedging motive cannot account for. In particular, we show that cross-strike price formation dynamics explain the “volatility smile”—a well-documented empirical regularity in option-implied volatilities that existing reduced-form models derived from Black-Scholes do not fully capture.111 Numerous reduced-form models have attempted to capture the volatility smile by modifying components within Black-Scholes-type frameworks. Prominent approaches include the jump-diffusion models (Merton (1976), Bates (1991)), stochastic volatility models (Heston (1993), Bates (1996), Duffie et al. (2000), Britten-Jones and Neuberger (2000), Bates (2000), Aït-Sahalia and Kimmel (2007), Christoffersen et al. (2010), Aït-Sahalia et al. (2021)), and local volatility models (Dupire (1994), Berestycki et al. (2002), and Carr and Cousot (2012)).
@@
Empirical research has consistently shown that trades and prices across options markets reflect information beyond the expected payoff. Pan and Poteshman (2006) and Ni et al. (2008) find that option trading volumes predict future stock returns and volatility, while Goyal and Saretto (2009) show that volatility mispricing can be exploited through option straddles. Other studies highlight how option order flows can signal firm value and market expectations of future volatility—see Roll et al. (2009), Augustin et al. (2019), and Cao et al. (2005). Despite substantial empirical evidence of joint price formation across option markets, existing theoretical models remain limited to no-option versus single-option comparisons (e.g., Back (1993); Biais and Hillion (1994); Easley et al. (1998); Collin-Dufresne et al. (2021)). This leaves an empirical-theoretical gap in understanding how trades and prices across options jointly reflect broader information. Our model bridges this gap by directly capturing the richness of information embedded across option trades and prices. It provides a unified theoretical framework that consolidates empirical findings and suggests new avenues of empirical research.
@@
Different from our strategic setting, there have also been theoretical studies on non-strategic informed trading across multiple securities, such as Admati (1985), Malamud (2015), and Chabakauri et al. (2022). The non-strategic models cannot explain the common derivatives trading strategies or observed cross-strike empirical regularities. Moreover, Admati (1985) and Chabakauri et al. (2022) impose restrictive parametric assumptions on the payoff distribution, and Malamud (2015) depends on properties intrinsic to the continuum. We make no parametric assumptions on the payoff, and our results do not depend on the discrete or continuum modeling choice. Additionally, we derive testable predictions linking the underlying payoff characteristics, such as higher moments, to observed price impacts across options. This new insight is unique to the strategic setting.
@@
The rest of the paper proceeds as follows. Section 2 demonstrates the basic intuition of our results through a simplified setting. Section 3 presents the general model. Section 4 identifies the intuitive sufficient statistic for the uninformed trader’s cross-market inferences. Section 5 analyzes the informed trader’s cross-market price impact. Section 6 establishes equilibrium through a canonical reduction of the general trading game. This canonical reduction is a key modeling innovation that yields a simple equilibrium for our general model, characterized by a single endogenous constant. For those readers not wishing to delve into the precise details, the broad intuition of the resulting equilibrium extends directly from Section 2 and can be gleaned from Example 6.7. Section 7 discusses the characterization of price discovery across contingent claims: the informed demand, price impact, and information efficiency of prices. Section 8 explores the immediate empirical predictions of our model for price discovery across option markets. Section 9 concludes. Formal assumptions and proofs are provided in the Appendix.
@@ 3 General Model
Suppose the states $x_{1}<x_{2}<\cdots$ are the possible $t=1$ prices of an underlying asset. In this case, any portfolio $W(\,\cdot\,)$ of AD securities can be replicated using a combination of the underlying asset and options: [^10] See Back (2010) Exercise 3.5. Let $x_{i_{0}}$ be the expected $t=1$ asset price under the market maker’s prior at $t=0$. For any $W(\,\cdot\,)$, we can find $(a_{i})$ such that
@@
Options in Kyle (1985)
In the options formulation, the single-asset case translates as follows.
@@
(i) Suppose the payoff distributions specified by the insider’s signal differ only in their expected values, and there are no option markets. In this case, our model reduces to a Kyle (1985) model in which only the underlying asset is traded.
(ii) Conversely, if the expected payoff remains identical across possible payoff distributions, then the insider’s demand for the underlying asset is zero. In other words, exploiting information beyond the expected payoff requires options.
Proposition 3.3(i) applies only when there are no option markets. In other words, the linear equilibrium (5) of Kyle (1985) is not robust with respect to the introduction of options. Options allow the insider to reduce demand optimally for lower-probability strikes, which makes his demand necessarily nonlinear. See Example 7.1 for an illustration.
@@ 4.2 Sufficient Statistics
In Theorem 4.1, we extend this intuition to arbitrary information structures. For each signal $s_{k}$, the market maker projects the order flow $\omega$ onto the demand he expects from the insider conditional on $s_{k}$, $\widetilde{W}(\,\cdot\,|s_{k})$. His posterior probability of signal $s_{k}$ is determined by the resulting noise-adjusted projection coefficient:
@@ 6.1 Canonical Game
The canonical transformation (24) normalizes the trading game by the insider’s information intensity. For intuition, suppose the $\eta(\,\cdot\,|s_{k})$’s have non-overlapping payoffs across states, such as in Section 2. In this case, there is zero covariation of payoffs across signals, which results in a diagonal information intensity matrix ${\bf L}$. The $k$-th diagonal entry of ${\bf L}$ is $c_{k}^{2}=\sum\limits_{i}\frac{\eta(x_{i}|s_{k})^{2}}{\sigma_{i}^{2}}$, which is the noise-adjusted payoff variation across securities conditional on $s_{k}$. The canonical transformation rescales each $\eta(\cdot|s_{k})$ by $\frac{1}{c_{k}}$, i.e., it normalizes the payoff variations to one across signals, thereby transforming ${\bf L}$ to the identity matrix. [^15] We note again that our proofs apply to any positive function $\eta(\,\cdot\,|\,\cdot\,)\colon X\times S\rightarrow\mathbb{R}$, including the rescaled payoffs. This normalization yields an isomorphic game that is invariant with respect to $\eta(\,\cdot\,|\,\cdot\,)$ and $(\sigma_{i})$ and symmetric across signals.
@@ 6.2 Symmetric Equilibrium
where ${\bf Q}=\mathbf{I}-\frac{1}{K}\bar{e}\bar{e}^{T}\in\mathbb{R}^{K\times K}$, with $\mathbf{I}$ denoting the identity matrix and $\bar{e}$ the vector of $1$’s.
@@ 7.1 Informed Demand
$W^{*}(\,\cdot\,|s_{k})=\sum_{l=1}^{K}\beta^{(k)}_{l}\eta(\,\cdot\,|s_{l}),\;\;\mbox{where $\beta^{(k)}$, the $k$-th column of $\alpha^{*}{\bf L}^{-1}{\bf Q}$, defines the allocation weights.}$. (35)
Applying the replication formula (7) yields the equivalent options portfolio.
@@
  • Step 1: Form the initial portfolio by buying $1-\frac{1}{K}$ shares of the observed distribution $\eta(\,\cdot\,|s_{l})$ and selling $\frac{1}{K}$ shares of each remaining distribution $\eta(\,\cdot\,|s_{l})$, $l\neq k$.
@@
Step 1 forms the initial portfolio according to the allocation weights specified by the $k$-th column of ${\bf Q}$. Step 2 adjusts the allocations to account for the insider’s information intensity ${\bf L}$; higher information intensity forces the insider to reduce the size of his trades across securities and signals. Step 3 scales the portfolio by $\alpha^{*}$ to optimize with respect to the market maker’s prior uncertainty over signals (see again Figure 2).
@@
The AD informed demand (35) extends the intuition behind the single-asset informed demand (5) to the contingent claims setting. This is made clear by comparing the AD informed demand in its canonical form (defined in Corollary 6.3) with the single-asset informed demand—both being determined by the information intensity and signal effect:

### v36

@@ 1 Introduction
A fundamental economic function of prices is to convey information. As informed agents trade to exploit their private knowledge about an asset’s expected payoff, that information becomes incorporated into the asset price. More broadly, agents can have information about an asset’s payoff beyond its mean—such as volatility, skew, or tail risk—that can only be exploited through trading derivatives. This utility of derivatives for targeting various aspects of the payoff distribution has been well understood since Ross (1976). These derivative trading strategies are now standard practice, as evidenced by widely used textbooks such as Hull (2015). This means derivative prices incorporate a much richer set of information than just the asset’s expected payoff.
Yet, the existing literature lacks a basic model articulating how derivative markets systematically incorporate diverse information signals into prices. This theoretical gap reflects the modeling challenges posed by the nonlinearity of derivative payoffs, the interdependence of trades and prices, and the resulting complex information dynamics across these markets. We address this gap by unifying two seminal frameworks: Arrow and Debreu (1954), which underpins contingent claim pricing, and Kyle (1985), a canonical model of informed trading. Our model delivers general characterizations of the informed demand, price impact, and information efficiency of prices across derivative markets.
In our model, an informed trader has private information about the probabilities of future states and trades state-contingent claims. In other words, the informed trader has information about the payoffs of Arrow-Debreu (AD) securities and trades in the AD markets. We impose no restriction on the probability distribution across states or the type of information available to the informed trader.
Replicating derivatives—such as options—using AD securities yields specializations of our model that account for the structural linkages across derivative instruments. In an options formulation, the informed trader has arbitrary private information about an underlying asset’s payoff distribution and trades options.
@@
The joint supply schedules for all AD securities constitute a pricing kernel that prices all derivatives under asymmetric information. While derivative pricing has been extensively studied from the complete-information, hedging perspective since Black and Scholes (1973), we show that informed trading better explains observed price behavior that hedging motives do not account for. In particular, our pricing kernel endogenously generates the “volatility smile”—a well-documented empirical regularity of option-implied volatilities. By contrast, reduced-form models descended from Black-Scholes have attempted to fit the smile ex post, but they can neither generate nor explain it.111 Numerous reduced-form models have attempted to fit the volatility smile by modifying components within Black-Scholes-type frameworks. Prominent approaches include the jump-diffusion models (Merton (1976), Bates (1991)), stochastic volatility models (Heston (1993), Bates (1996), Duffie et al. (2000), Britten-Jones and Neuberger (2000), Bates (2000), Aït-Sahalia and Kimmel (2007), Christoffersen et al. (2010), Aït-Sahalia et al. (2021)), and local volatility models (Dupire (1994), Berestycki et al. (2002), and Carr and Cousot (2012)).
@@
A large empirical literature shows that trades and prices across options markets reflect information beyond just the expected payoff. Pan and Poteshman (2006) and Ni et al. (2008) find that option trading volumes predict stock returns and volatility, while Goyal and Saretto (2009) document profitable straddle trades that exploit volatility mispricing. Additional evidence confirms that option order flows contain firm-value and tail-risk information (Roll et al. (2009), Augustin et al. (2019), and Cao et al. (2005)). Despite robust empirical evidence of joint price formation across option markets, existing theoretical models remain confined to no-option versus single-option comparisons (e.g., Back (1993); Biais and Hillion (1994); Easley et al. (1998); Collin-Dufresne et al. (2021)). Our model bridges this empirical-theoretical gap. It consolidates empirical findings and generates new testable cross-option predictions. For example, while Pan and Poteshman (2006) show that signed option volume contains directional information, we predict that cross-strike volume imbalance is the precise carrier transmitting that signal. We also offer a theoretical explanation for why option straddle demand predicts realized volatility, consistent with Ni et al. (2008). Section 8 details these and additional testable implications.
@@
In contrast to our strategic setting, there have been prior theoretical studies examining non-strategic informed trading across multiple securities, such as Admati (1985), Malamud (2015), and Chabakauri et al. (2022). The non-strategic models cannot explain the common derivatives trading strategies or observed cross-strike empirical regularities. Moreover, Admati (1985) and Chabakauri et al. (2022) impose restrictive parametric assumptions on the payoff distribution, while Malamud (2015) relies on properties intrinsic to the continuum. We make no parametric assumptions on the payoff, and our results do not depend on the discrete or continuum modeling choice. Additionally, we derive empirical predictions linking characteristics of the underlying payoff to observed cross-strike price impact. This new insight is unique to the strategic setting.
@@
The remainder of the paper is organized as follows: Section 2 lays out the core intuition of our results in a simplified setting. Section 3 presents the general model. Section 4 derives the sufficient statistic governing the uninformed trader’s cross-market inferences. Section 5 characterizes the informed trader’s cross-market price impact. Section 6 establishes equilibrium via a canonical reduction of the general trading game. This canonical reduction is a key modeling innovation that yields a simple equilibrium in our general model, fully characterized by a single endogenous constant. For readers not wishing to delve into details, the broad intuition behind equilibrium can be gleaned from Example 6.7, which directly connects to the simplified setting of Section 2. Section 7 examines price discovery across contingent claims: the informed demand, price impact, and information efficiency of prices. Section 8 explores the empirical predictions of our model for price discovery across option markets. Section 9 concludes. Formal assumptions and proofs are provided in the Appendix.
@@ 3 General Model
Suppose the states $x_{1}<x_{2}<\cdots$ are the possible $t=1$ prices of an underlying asset. In this case, any portfolio $W(\,\cdot\,)$ of AD securities can be replicated using a combination of the underlying asset and options: [^10] See Back (2010), Exercise 3.5. Let $x_{i_{0}}$ be the expected $t=1$ asset price under the market maker’s prior at $t=0$. For any $W(\,\cdot\,)$, we can find $(a_{i})$ such that
@@
Options in the Kyle (1985) Framework
The options formulation relates to the single-asset case as follows:
@@
(i) If the payoff distributions differ across signals only in their expected values, and there are no option markets, then our model reduces to the Kyle (1985) framework, where only the underlying asset is traded.
(ii) Conversely, if all possible payoff distributions have the same expected value, then the insider has no incentive to trade the underlying asset. In other words, exploiting information beyond the expected payoff requires options.
Proposition 3.3(i) applies only when there are no option markets. In other words, the linear equilibrium (5) in Kyle (1985) is not robust to the introduction of options. Options allow the insider to optimally scale back his demands for low-probability strikes, resulting in an overall demand that is inherently nonlinear. This point will be illustrated explicitly in Example 7.1 below.
@@ 4.2 Sufficient Statistics
In Theorem 4.1, we extend the above intuition to arbitrary information structures. For each signal $s_{k}$, the market maker projects the order flow $\omega$ onto the demand he expects from the insider conditional on $s_{k}$, $\widetilde{W}(\,\cdot\,|s_{k})$. His posterior probability of signal $s_{k}$ is determined by the resulting noise-adjusted projection coefficient:
@@ 6.1 Canonical Game
The canonical transformation (24) normalizes the trading game by the insider’s information intensity. To see this intuition clearly, return to the simplified scenario of Section 2 where the payoff distributions $\eta(\,\cdot\,|s_{k})$ do not overlap across states. In this scenario, because there is zero covariation of payoffs across signals, the information intensity matrix ${\bf L}$ is diagonal. The $k$-th diagonal entry of ${\bf L}$ is $c_{k}^{2}=\sum\limits_{i}\frac{\eta(x_{i}|s_{k})^{2}}{\sigma_{i}^{2}}$, which is the noise-adjusted payoff variation across securities conditional on $s_{k}$. The canonical transformation rescales each payoff distribution $\eta(\,\cdot\,|s_{k})$ by the factor $\frac{1}{c_{k}}$, thus converting ${\bf L}$ to the identity matrix. In other words, this transformation normalizes the payoff variations across signals to unity. [^15] We note again that our proofs apply to any positive function $\eta(\,\cdot\,|\,\cdot\,)\colon X\times S\rightarrow\mathbb{R}$, including the rescaled payoffs considered here. This normalization yields an isomorphic game that is invariant with respect to $\eta(\,\cdot\,|\,\cdot\,)$ and $(\sigma_{i})$ and symmetric across signals.
@@ 6.2 Symmetric Equilibrium
where
[…]
@@ 7.1 Informed Demand
[…]
Applying the replication formula (7) produces the equivalent options portfolio.
@@
  • Step 1: Form the initial portfolio by buying $1-\frac{1}{K}$ shares of the observed distribution $\eta(\,\cdot\,|s_{k})$ and selling $\frac{1}{K}$ shares of each of the other distributions $\eta(\,\cdot\,|s_{l})$, $l\neq k$.
@@
Step 1 forms the initial long-short portfolio based on the observed signal. Step 2 adjusts the allocations to account for the insider’s information intensity; greater information intensity compels the insider to scale down his trades across securities and signals. Step 3 scales the portfolio by the endogenous constant $\alpha^{*}$ to optimize with respect to the market maker’s prior uncertainty over signals (see again Figure 2).
@@
The AD informed demand (36) extends the intuition behind the single-asset informed demand (5) to contingent claims. This parallel becomes clear when comparing the AD informed demand in its canonical form (defined in Corollary 6.3) with the single-asset informed demand, as both are determined by the information intensity and signal effect: